OpenStudy (anonymous):

Factor the four-term polynomial. mn - 15 + 3m - 5n A. (m + 3)(n - 5) B. (n - 3)(m - 5) C. (n + 3)(m - 5)

OpenStudy (uri):

mn+3m - 15-5n = m(n + 3) - 5(n+3) = (n+3) (m-5) C is the answer.

OpenStudy (anonymous):

ok thank you